Suppose that an economy is composed of three classes. Labourers, traders and capital owners. The size of each of the classes are 60%, 30% and 10% respectively. Laborers produce all the output in this economy by using capital borrowed from the capitalists and then use the services of traders to market their produce. For this, the labourers have to pay 40% and 20% of their produce to the capitalists and traders respectively. a. Draw the Lorenz curve for the income distribution of this economy. Calculate the Gini Coefficient. b. Suppose that the capitalist are taxed 10% of their income and this is distributed as transfers to the labourers. Draw the new Lorenz curve and compare income inequality between the market income distribution and the disposable income distribution


1
Expert's answer
2021-11-14T17:59:21-0500

(a)

Lorenz curve for the income distribution of the economy:



The Gini coefficient:

"=\frac{A}{A+B}"

"=1-(2\times0.2)\times(0.133+0.267+0.20+0.40)"

"=0.6\times1=0.6"

(b)

The new Lorenz curve:



The income inequality that exists between the market income distribution and disposable income distribution will slightly improve towards level of income equality.